OpenStudy (anonymous):

Let f(x) = x + 5 and g(x) = x2. Find (g ͦ f)(6)? a. 81 b. 36 c. 121 Can you show your work too? I've asked the teacher but I still don't get it and I need to know how. I've tried every math site and my textbook, with the math I tried to do I think it might be 36 but I'm not sure...please help, please I really need the help.

OpenStudy (freckles):

so (g o f)(6) means g(f(6)) this means we need to start with the inside f(6) what is f(6) well f(x)=x+5 so f(6)=6+5=?

OpenStudy (freckles):

after we find f(6) we will need to find g(f(6))

OpenStudy (freckles):

@Vendela are you there

OpenStudy (anonymous):

Yeah I'm listening

OpenStudy (freckles):

well can you tell me what f(6) is?

OpenStudy (anonymous):

umm...is it 11,because of the 6+5?

OpenStudy (freckles):

yes so we have g(f(6))=g(11)

OpenStudy (freckles):

but g(x)=x^2

OpenStudy (freckles):

so g(11)=?

OpenStudy (anonymous):

66?

OpenStudy (freckles):

11^2 isn't 66

OpenStudy (anonymous):

oh whoops wrong number lol okay it'd be 121

OpenStudy (freckles):

\[(g ^o f)(6)=g(f(6))=g(6+5)=g(11)=11^2=121 \text{ yes! }\]

OpenStudy (anonymous):

OMG thank you soooooooooo much!!!!!!!1

OpenStudy (freckles):

do you want to try to find (f o g)(6)?

OpenStudy (anonymous):

*!

OpenStudy (anonymous):

oh yes

OpenStudy (freckles):

so that means f(g(6)) right so this time g(6) is inner most thing to evaluate

OpenStudy (freckles):

g(6) tells us we are going to use the function called g and g(x)=x^2 so g(6)=?

OpenStudy (anonymous):

36

OpenStudy (freckles):

so we have f(g(6))=f(36)

OpenStudy (freckles):

do you know how to evaluate f(36)?

OpenStudy (anonymous):

no not really

OpenStudy (freckles):

well isn't f given by f(x)=x+5

OpenStudy (anonymous):

okayy

OpenStudy (freckles):

since f(x)=x+5 then f(36)=36+5

OpenStudy (anonymous):

41

OpenStudy (freckles):

\[(g ^o f)(6)=g(f(6))=g(6+5)=g(11)=11^2=121 \\ (f ^o g)(6)=f(g(6))=f(6^2)=f(36)=36+5=41\]

OpenStudy (freckles):

now i hope you notice that was just an extra question

OpenStudy (anonymous):

so the answer is 121?

OpenStudy (freckles):

just for more practice

OpenStudy (freckles):

yes

OpenStudy (freckles):

you might want to practice these compositions more

OpenStudy (anonymous):

Thank you so so so so much.You are honestly better than my math teacher:P how do I medal+fan?

OpenStudy (freckles):

you can medal and fan if you want but it is not required the best response button on this post will give me a medal scrolling over my name will show you a tab to become a fan

OpenStudy (anonymous):

awesome!! thank you again!!

OpenStudy (freckles):

np